# Transfer Learning Between U.S. Presidential Elections: How Should We Learn From A 2020 Ad Campaign To Inform 2024 Ad Campaigns?

### v1

@@ 2.3 Causal Identification
Assumption 2.1 is automatically satisfied if the source data is from a randomized controlled trial, such as our source data from Aggarwal et al., (2023). Also, to identify the TATE, SUTVA is not necessarily for the target population (i.e. $S_{i}=0$). This is because the identification strategy under transfer learning is based on transferring information about the potential outcomes, not the observed outcomes.
@@
The first part of Assumption 2.2 will be violated if there are some values of the shared covariates ${\bf V}_{i}$ that are only observed in the target population, for instance if Aggarwal et al., (2023) focused only on young voters and the target population consists of voters from all ages. The second part of Assumption 2.2 will be violated if the target sample size is much smaller than the source sample size. Because both parts depend solely on observable quantities, Assumption 2.2 can be checked with data; see Figure 6.1 for an example.
Assumption 2.3, henceforth referred to as transportability, states that conditional on the shared covariates ${\bf V}_{i}$, the distributions of the potential outcomes are identically distributed between the source and the target populations. The assumption is violated if the distribution of the potential outcomes differ between the source and the target populations after adjusting for ${\bf V}_{i}$. For example, if ${\bf V}_{i}$ only contains political party, Assumption 2.3 will be violated if within each political party, voter turnout under treatment or control is different between the 2020 and 2024 elections. Unfortunately, unlike Assumption 2.2, Assumption 2.3 depends on counterfactual quantities and cannot be checked with data. Furthermore, unlike strong ignorability in Assumption 2.1, we are not aware of a feasible experimental design to guarantee Assumption 2.3 in electoral contexts222 One experimental design to satisfy Assumption 2.3 is for the investigator to randomize the selection of study units into the source sample (Tipton,, 2013; Cook,, 2014; Tipton and Peck,, 2017). In our data analysis, this design implies Aggarwal et al., (2023) randomized voters to be either in their 2020 study or to be a registered voter in PA for the 2024 election. We believe that this design is impractical and the “nonignorable selection bias” into the target or the source sample is unavoidable. . This is the main motivation for us to embed sensitivity analysis within transfer learning so that our framework does not reply on Assumption 2.3; see Section 3.
@@
In words, $\theta$ is identified by first averaging the conditional average treatment effect in the source population (i.e., CATE or formally, $\mu_{1}({\bf X}_{i})-\mu_{0}({\bf X}_{i})$) over the shared covariates ${\bf V}_{i}$ (i.e., the inner expectation in equation (2.1)) and second, averaging this quantity among units in the target population (i.e., the outer expectation in equation (2.1)). For estimation of $\theta$ under Assumptions 2.1-2.3, see Zeng et al., (2023).
@@ 3 Sensitivity Analysis of Transportability
As discussed above, suppose transportability (i.e., Assumption 2.3) no longer holds and we measure the departure from it by the sensitivity parameter $\gamma_{a}\in\mathbb{R}$ for $a\in\{0,1\}$. The parameter $\gamma_{a}$ is the logarithm of the odds ratio of counterfactual outcomes between the target and source populations:
$\gamma_{a}$ $={\textnormal{log}}\left(\frac{{\rm ODD}_{a}({\bf v},0)}{{\rm ODD}_{a}({\bf v},1)}\right),\ {\rm ODD}_{a}({\bf v},s)=\frac{\mathbb{P}(Y^{(a)}_{i}=1\mid{\bf V}_{i}={\bf v},S_{i}=s)}{\mathbb{P}(Y^{(a)}_{i}=0\mid{\bf V}_{i}={\bf v},S_{i}=s)},\ s\in\{0,1\},{\bf v}\in{\cal V}.$ (3.1)
When $\gamma_{a}=0$, i.e., the distribution of $Y^{(a)}_{i}$ is identical in the source and target populations conditioned on the shared covariates ${\bf V}_{i}$, Assumption 2.3 holds, and we recover the setting in Zeng et al., (2023). As $\gamma_{a}$ moves away from 00, the degree of violation of transportability increases. For example, in our data analysis, $\gamma_{1}=0.05$ means that the counterfactual odd of voting in 2024 is $\exp(0.05)\approx 1.05$ times that in 2020 when a registered voter, possibly contrary to fact, gets negative ads against Trump. The sensitivity parameter $\gamma_{a}$ is not identified and cannot be estimated. In this and the subsequent sections, we discuss the identification and estimation of the TATE with a given $\gamma_{a}$, and then in Section 5 we discuss a data-driven approach to calibrate this parameter.
Under the sensitivity model (3.1) and for a given $\gamma_{a}\in\mathbb{R}$, the expected counterfactual outcome under treatment level $a\in\{0,1\}$ can be identified as follows.
@@
Suppose Assumptions 2.1, 2.2 and the sensitivity model in model (3.1) hold. For a given $\gamma_{a}\in\mathbb{R}$, the expected counterfactual outcome under treatment level $a\in\{0,1\}$ is
$\mathbb{E}[Y^{(a)}_{i}\mid S_{i}=0]$ $=\mathbb{E}\left[\dfrac{\exp(\gamma_{a})\rho_{a}({\bf V}_{i})}{\exp(\gamma_{a})\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})}\Bigg{|}S_{i}=0\right]$ (3.2)
$=\theta_{a}(\gamma_{a}).$
@@
To highlight the inclusion of the sensitivity analysis, we denote the mean counterfactual outcome under treatment level $a$ by $\theta_{a}(\gamma_{a})$ and the TATE by $\theta_{1}(\gamma_{1})-\theta_{0}(\gamma_{0})$. Despite the expanded notation, the interpretation of $\theta_{a}(\gamma_{a})$ as an average of the counterfactual outcome $Y^{(a)}_{i}$ in the target population remains the same regardless of the value of $\gamma_{a}$. For example, in our data analysis, if $\gamma_{1}=0$, $\theta_{1}(0)$ is the proportion of registered PA voters who would vote in the 2024 election if all voters were assigned to anti-Trump digital ads and transportability held. If $\gamma_{1}=0.1$, $\theta_{1}(0.1)$ is the proportion of registered PA voters who would vote in the 2024 election if all voters were assigned to anti-Trump digital ads and transportatbility was violated by $\gamma_{1}=0.1$.
@@
We conclude this section with a couple of remarks on the sensitivity model (3.1). First, this model was first proposed by Robins et al., (2000) as a non-parametric (just) identified model for describing selection bias in missing data. The model was later called an exponential tilting model (Rotnitzky et al.,, 2001; Birmingham et al.,, 2003), an extrapolation-factorization model (Linero and Daniels,, 2018) and Tukey’s representation (Franks et al.,, 2020). The model was also used to conduct sensitivity analysis for unmeasured confounding in causal inference (Franks et al.,, 2020; Scharfstein et al.,, 2021) and for violation of the transportability assumption in generalizability (Dahabreh et al.,, 2022). In particular, when ${\cal V}={\cal X}$, Lemma 3.1 recovers the identification of the TATE in Dahabreh et al., (2022). Second, we choose this model for sensitivity analysis as it (a) posits no testable implications on the data, (b) makes statistical inference tractable (e.g., asymptotic normality), and (c) has a simple, odds ratio interpretation. Third, the sensitivity model can be extended in various ways. For example, it can be extended to handle a continuous, counterfactual outcome where the extension would tilt the entire density of the counterfactual outcome; see Section A of the Appendix where we discuss identification, estimation, and interpretation of the TATE under a sensitivity model for a continuous, counterfactual outcome. Also, extensions to model (3.1) can induce unmeasured differences as a function of the outcome $Y^{(a)}_{i}$ and covariates ${\bf V}_{i}$; see Franks et al., (2020) and Scharfstein et al., (2021) for examples. In addition, model (3.1) can be reformulated under a selection model (see Section A of the Appendix) or under an $R^{2}$-based model (Franks et al.,, 2020).
@@ 4.1 Outcome Regression and Percentile Bootstrap
The analysis pipeline in this section is most appropriate when ${\cal X}$ is discrete or, more generally, when the outcome regression model $\rho_{a}({\bf v})$ can be consistently estimated at a parametric rate. This is the case in our data analysis where voter’s demographics are discrete variables. Even if ${\cal X}$ is not discrete, we suggest investigators begin with the analysis in this section since it is not only simple, but also the alternative analysis based on the efficient influence function (EIF) does not have doubly robust rates; see Section 4.2.
From equation (3.2), a natural estimator of $\theta_{a}(\gamma_{a})$ would be a plug-in estimator that takes a weighted average of an estimator of $\rho_{a}({\bf v})$, denoted as $\widehat{\rho}_{a}({\bf v})$, among the target sample:
$\widehat{\theta}_{\textnormal{OR,a}}(\gamma_{a})=\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\exp(\gamma_{a})\widehat{\rho}_{a}({\bf V}_{i})}{\exp(\gamma_{a})\widehat{\rho}_{a}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})}.$ (4.1)
Also, a simple estimator of $\widehat{\rho}_{a}({\bf v})$ is motivated by its definition where we regress $\widehat{\mu}_{a}({\bf x})$ on ${\bf v}$ using least squares and $\widehat{\mu}_{a}({\bf x})$ is an estimate of $\mu_{a}({\bf x})$. For example, if ${\cal X}$ is discrete, both $\widehat{\mu}_{a}({\bf x})$ and $\widehat{\rho}_{a}({\bf v})$ can be expressed as
$\widehat{\mu}_{a}({\bf x})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}Y_{i}\mathbf{1}(A_{i}=a,{\bf X}_{i}={\bf x})}{\sum_{i\in{\cal I}_{s}}\mathbf{1}(A_{i}=a,{\bf X}_{i}={\bf x})},\ \widehat{\rho}_{a}({\bf v})=\dfrac{\sum_{i\in{\cal I}_{s}}\widehat{\mu}_{a}({\bf X}_{i})\mathbf{1}({\bf V}_{i}={\bf v})}{\sum_{i\in{\cal I}_{s}}\mathbf{1}({\bf V}_{i}={\bf v})},\ {\bf x}\in{\cal X},{\bf v}\in{\cal V},$ (4.2)
where $\mathbf{1}(\cdot)$ is the indicator function. In the discrete case, the estimators in equation (4.2) are consistent. For inference, we recommend a nonparametric, percentile bootstrap (Efron,, 1979) where the source and the target data are resampled separately and we take the $\alpha/2$ and $21-\alpha/$ 2 quantiles of the bootstrapped estimates of $\widehat{\theta}_{\textnormal{OR,a}}(\gamma_{a})$, denoted $\widehat{L}_{a}(\gamma_{a};1-\alpha)$ and $\widehat{U}_{a}(\gamma_{a};1-\alpha)$ respectively. These quantiles are used to construct a $(1-\alpha)$ confidence interval (CI), denoted as $\widehat{\text{CI}}_{\textnormal{OR,a}}(\gamma_{a};1-\alpha)=\left[\widehat{L}_{a}(\gamma_{a};1-\alpha),\widehat{U}_{a}(\gamma_{a};1-\alpha)\right]$.
Formally, suppose $\rho_{a}({\bf v})$ is indexed by a finite-dimensional parameter $\boldsymbol{\eta}_{a}$ and denote its estimate as $\rho_{a}({\bf v},\widehat{\boldsymbol{\eta}}_{a})$. Theorem 4.1 shows that under regularity conditions, the plug-in estimator $\widehat{\theta}_{\textnormal{OR,a}}(\gamma_{a})$ in (4.1) is consistent and the nonparametric, percentile bootstrap leads to a valid CI.
@@ 4.2 Efficient Influence Function
The analysis pipeline in this section is based on the efficient influence function (EIF) and is more general than that in Section 4.1, especially when ${\cal V}$ is not discrete and the propensity score is unknown. However, it is more complex and requires estimating multiple nuisance functions. To motivate the estimator, we first derive the EIF of $\theta_{a}(\gamma_{a})$ in Theorem 4.2.
@@
$\textnormal{EIF}({\bf O}_{i},\theta_{a}(\gamma_{a}))=$ $\dfrac{S_{i}w({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\exp(\gamma_{a})}{[\exp(\gamma_{a})\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\pi({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\pi({\bf X}_{i})}\right\}\{Y_{i}-\mu_{a}({\bf X}_{i})\}\right.$
$+\mu_{a}({\bf X}_{i})-\rho_{a}({\bf V}_{i})\bigg{]}+\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\left[\dfrac{\exp(\gamma_{a})\rho_{a}({\bf V}_{i})}{\exp(\gamma_{a})\rho_{a}({\bf V})+1-\rho_{a}({\bf V}_{i})}-\theta_{a}(\gamma_{a})\right],$
@@
The EIF involves four nuisance quantities: the propensity score $\pi({\bf X}_{i})$, the ratio of densities of the covariates between the target and the source population $w({\bf V}_{i})$, and two outcome regression functions $\mu_{a}({\bf X}_{i})$ and $\rho_{a}({\bf V}_{i})$. When transportability holds, i.e., $\exp(\gamma_{a})=1$, Theorem 4.2 reduces to the EIF in Zeng et al., (2023).
We follow the modern trend in causal inference where we use cross-fitting (Chernozhukov et al.,, 2017; Kennedy,, 2022) and the EIF to estimate $\theta_{a}(\gamma_{a})$. Specifically, we randomly partition the source and target sample indices ${\cal I}_{s}$ and ${\cal I}_{t}$ into $K$ disjoint sets, ${\cal I}_{s}^{(k)}$ and ${\cal I}_{t}^{(k)}$, respectively, for $k=1,2,\cdots,K$, and let ${\cal I}^{(k)}={\cal I}_{s}^{(k)}\cup{\cal I}_{t}^{(k)}$. For each $k$, the nuisance functions are estimated with data in ${\cal I}\setminus{\cal I}^{(k)}$ and they are denoted as $\widehat{\pi}^{(k)}({\bf x})$, $\widehat{\mu}_{a}^{(k)}({\bf x})$, $\widehat{w}^{(k)}({\bf v})$ and $\widehat{\rho}_{a}^{(k)}({\bf v})$. We then plug them into the “uncentered” EIF and evaluate it with the data in ${\cal I}^{(k)}$, i.e.,
$\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\gamma_{a})=$ $\dfrac{1}{|{\cal I}_{s}^{(k)}|}\sum_{i\in{\cal I}_{s}^{(k)}}\dfrac{\exp(\gamma_{a})\widehat{w}^{(k)}({\bf V}_{i})}{[\exp(\gamma_{a})\widehat{\rho}_{a}^{(k)}({\bf V}_{i})+1-\widehat{\rho}_{a}^{(k)}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\widehat{\pi}^{(k)}({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\widehat{\pi}^{(k)}({\bf X}_{i})}\right\}\{Y_{i}-\widehat{\mu}_{a}^{(k)}({\bf X}_{i})\}\right.$
$\left.\quad{}\quad{}\quad{}\quad{}+\widehat{\mu}_{a}^{(k)}({\bf X}_{i})-\widehat{\rho}_{a}^{(k)}({\bf V}_{i})\right]+\dfrac{1}{|{\cal I}_{t}^{(k)}|}\sum_{i\in{\cal I}_{t}^{(k)}}\dfrac{\exp(\gamma_{a})\widehat{\rho}_{a}^{(k)}({\bf V}_{i})}{\exp(\gamma_{a})\widehat{\rho}_{a}^{(k)}({\bf V}_{i})+1-\widehat{\rho}_{a}^{(k)}({\bf V}_{i})}.$
Finally, we take an average of $\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\gamma_{a})$ to arrive at the EIF-based cross-fitting estimator of $\theta_{a}(\gamma_{a})$, which we denote as $\widehat{\theta}_{\textnormal{EIF,a}}(\gamma_{a})=K^{-1}\sum_{k=1}^{K}\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\gamma_{a})$. Theorem 4.3 shows that $\widehat{\theta}_{\textnormal{EIF,a}}(\gamma_{a})$ is consistent, asymptotically normal, and semiparametrically efficient.

### v2

@@ 2.3 Causal identification
Assumption 2.1 is automatically satisfied if the source data is from a randomized controlled trial, such as our source data from Aggarwal et al., (2023). Also, to identify the TATE, SUTVA is not necessarily for the target population (i.e. $S_{i}=0$). This is because identification is based on transferring information about the potential outcomes, not the observed outcomes.
@@
The first part of Assumption 2.2 will be violated if there are some values of the shared covariates ${\bf V}_{i}$ that are only observed in the target population, for instance if Aggarwal et al., (2023) focused only on young voters and the target population consists of voters from all ages. The second part of Assumption 2.2 excludes the case where the target sample size is much smaller than the source sample size. Because both parts depend solely on observable quantities, Assumption 2.2 can be checked with data; see Figure 6.1 for an example.
Assumption 2.3, referred to as transportability, states that conditional on the shared covariates ${\bf V}_{i}$, the distributions of the potential outcomes are identical between the source and the target populations. The assumption is violated if the distribution of the potential outcomes differ between the source and the target populations after adjusting for ${\bf V}_{i}$. For example, if ${\bf V}_{i}$ only contains political party, Assumption 2.3 will be violated if within each political party, voter turnout under treatment or control is different between the 2020 and 2024 elections. Unfortunately, unlike Assumption 2.2, Assumption 2.3 depends on counterfactual quantities and cannot be checked with data. Furthermore, unlike strong ignorability in Assumption 2.1, we are not aware of a feasible experimental design to guarantee Assumption 2.3 in electoral contexts.222 A study design that satisfies Assumption 2.3 is to randomize the selection of study units into the source data (Tipton,, 2013; Tipton and Peck,, 2017). In our data analysis, this design implies Aggarwal et al., (2023) randomized voters to be either in their 2020 study or to be a registered voter in PA for the 2024 election and we believe that this design is impractical. This is the main motivation for us to embed sensitivity analysis within transfer learning so that our framework does not rely on Assumption 2.3.
@@
In words, $\theta$ is identified by first averaging the conditional average treatment (CATE) effect in the source population (i.e., $\mu_{1}({\bf X}_{i})-\mu_{0}({\bf X}_{i})$) over the shared covariates ${\bf V}_{i}$ (i.e., the inner expectation in equation (2.1)) and second, averaging this quantity among units in the target population (i.e., the outer expectation in equation (2.1)). For efficient and minimax estimation of $\theta$ under Assumptions 2.1-2.3, see Zeng et al., (2023).
@@ 3 Sensitivity analysis of transportability
As discussed above, suppose transportability (i.e., Assumption 2.3) no longer holds even after conditioning on $V_{i}$ and we measure the departure from it by the sensitivity parameter $\Gamma_{a}\in(0,\infty)=\mathbb{R}^{+}$ for $a\in\{0,1\}$. Specifically, the parameter $\Gamma_{a}$ is defined as the odds ratio of counterfactual outcomes between the target and source populations for a given ${\bf v}\in{\cal V}$:
$\Gamma_{a}$ $=\frac{{\rm ODD}_{a}({\bf v},0)}{{\rm ODD}_{a}({\bf v},1)},\ {\rm ODD}_{a}({\bf v},s)=\frac{\mathbb{P}(Y^{(a)}_{i}=1\mid{\bf V}_{i}={\bf v},S_{i}=s)}{\mathbb{P}(Y^{(a)}_{i}=0\mid{\bf V}_{i}={\bf v},S_{i}=s)},\ s\in\{0,1\},{\bf v}\in{\cal V}.$ (3.1)
When $\Gamma_{a}=1$, i.e., the conditional distributions of $Y^{(a)}_{i}$ given ${\bf V}_{i}$ are identical between the source and target populations (i.e., Assumption 2.3 holds). As $\Gamma_{a}$ moves away from $1$, the degree of violation of transportability increases. For example, in our data analysis, $\Gamma_{1}=1.05$ means that the counterfactual odd of voting in 2024 is $1.05$ times that in 2020 when a registered voter, possibly contrary to fact, gets negative ads against Trump. Similarly, $\Gamma_{1}=0.95$ means the counterfactual odd of voting in 2024 is $0.95$ times that in 2020 when a registered voter, possibly contrary to fact, gets negative ads about Trump.
Similar to other sensitivity analyses, the sensitivity parameter $\Gamma_{a}$ is not identifiable. Instead, investigators identify and estimate the TATE for a given $\Gamma_{a}$ and in doing so, study the sensitivity of the TATE when transportability is violated. Specifically, for a given $\Gamma_{a}\in\mathbb{R}^{+}$, the expected counterfactual outcome under treatment level $a\in\{0,1\}$ is identified as follows.
@@
Suppose Assumptions 2.1 and 2.2 hold. For a given $\Gamma_{a}\in\mathbb{R}^{+}$, the expected counterfactual outcome under treatment level $a\in\{0,1\}$ is
$\mathbb{E}[Y^{(a)}_{i}\mid S_{i}=0]$ $=\mathbb{E}\left[\dfrac{\Gamma_{a}\rho_{a}({\bf V}_{i})}{\Gamma_{a}\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})}\Bigg{|}S_{i}=0\right]=\theta_{a}(\Gamma_{a}).$ (3.2)
@@
To highlight the inclusion of the sensitivity analysis, we denote the mean counterfactual outcome under treatment level $a$ by $\theta_{a}(\Gamma_{a})$ and the TATE by $\theta_{1}(\Gamma_{1})-\theta_{0}(\Gamma_{0})$. Despite the expanded notation, the interpretation of $\theta_{a}(\Gamma_{a})$ as an average of the counterfactual outcome $Y^{(a)}_{i}$ in the target population remains the same regardless of the value of $\Gamma_{a}$. For example, in our data analysis, if $\Gamma_{1}=1$, $\theta_{1}(1)$ is the proportion of registered PA voters who would vote in the 2024 election if all voters were assigned to anti-Trump digital ads and transportability held. If $\Gamma_{1}=1.1$, $\theta_{1}(1.1)$ is the proportion of registered PA voters who would vote in the 2024 election if all voters were assigned to anti-Trump digital ads and transportatbility was violated by $\Gamma_{1}=1.1$.
@@
We conclude this section with a couple of remarks on the sensitivity model (3.1). First, this model was first proposed by Robins et al., (2000) as a non-parametric (just) identified model for describing selection bias in missing data. The model was later called an exponential tilting model (Rotnitzky et al.,, 2001; Birmingham et al.,, 2003) and an extrapolation-factorization model (Linero and Daniels,, 2018). The model was also used to conduct sensitivity analysis for unmeasured confounding in causal inference (Franks et al.,, 2020; Scharfstein et al.,, 2021) and for violation of the transportability assumption in generalizability (Dahabreh et al.,, 2022). In particular, when ${\cal V}={\cal X}$, Lemma 3.1 recovers the identification of the TATE in Dahabreh et al., (2022). Second, we choose this model for sensitivity analysis as it (a) posits no testable implications on the data, (b) makes statistical inference tractable (e.g., asymptotic normality), and (c) has a simple, odds ratio interpretation. Third, the sensitivity model can be extended in various ways. For example, it can be extended to handle a continuous, counterfactual outcome where the sensitivity model tilts the entire density of the counterfactual outcome; see Section A of the Appendix where we discuss identification, estimation, and interpretation of the TATE under a sensitivity model for a continuous, counterfactual outcome. Also, at the expense of more sensitivity parameters, model (3.1) can be extended to allow $\Gamma_{a}$ to depend on ${\bf V}_{i}$ and $Y^{(a)}_{i}$; see Franks et al., (2020) and Scharfstein et al., (2021) for examples. Fourth, model (3.1) can be reformulated under a selection model (see Section A of the Appendix) or under an $R^{2}$-based model (Franks et al.,, 2020).
@@ 4.1 Outcome regression and percentile bootstrap
The analysis pipeline in this section is appropriate when ${\cal X}$ is discrete or, more generally, when the outcome regression model $\rho_{a}({\bf v})$ can be consistently estimated at a parametric rate. This is the case in our data analysis where voter’s demographics are discrete variables. Even if ${\cal X}$ is not discrete, we suggest investigators begin with this analysis since it is not only simple, but also the alternative analysis based on the efficient influence function (EIF) is not doubly robust; see Section 4.2.
From the identification equation (3.2), a natural estimator of $\theta_{a}(\Gamma_{a})$ would be a plug-in estimator that takes a weighted average of an estimator of the outcome regression function $\rho_{a}({\bf v})$, denoted as $\widehat{\rho}_{a}({\bf v})$, among the target sample. We call this estimator the outcome regression (OR) estimator:
$\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})=\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})}{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})}.$ (4.1)
Also, from the definition of $\rho_{a}({\bf v})$ n Section 2.2, a simple estimator of $\widehat{\rho}_{a}({\bf v})$ is to regress $\widehat{\mu}_{a}({\bf x})$ on ${\bf v}$ using ordinary least squares (OLS) and $\widehat{\mu}_{a}({\bf x})$ is an estimate of $\mu_{a}({\bf x})$. If ${\cal X}$ is discrete, the OLS estimators of $\widehat{\mu}_{a}({\bf x})$ and $\widehat{\rho}_{a}({\bf v})$ can be expressed as
$\widehat{\mu}_{a}({\bf x})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}Y_{i}\mathds{1}(A_{i}=a,{\bf X}_{i}={\bf x})}{\sum_{i\in{\cal I}_{s}}\mathds{1}(A_{i}=a,{\bf X}_{i}={\bf x})},\ \widehat{\rho}_{a}({\bf v})=\dfrac{\sum_{i\in{\cal I}_{s}}\widehat{\mu}_{a}({\bf X}_{i})\mathds{1}({\bf V}_{i}={\bf v})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})},\ {\bf x}\in{\cal X},{\bf v}\in{\cal V},$ (4.2)
where $\mathds{1}(\cdot)$ is the indicator function. In the discrete case, the estimators in equation (4.2) are consistent. For a general discussion on estimating $\rho_{a}$, see Section 4.3.
For inference, we recommend a nonparametric, percentile bootstrap (Efron,, 1979) where the source and the target data are resampled separately and we take the $\alpha/2$ and $21-\alpha/$ 2 quantiles of the bootstrapped estimates of $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})$, denoted $\widehat{L}_{a}(\Gamma_{a};1-\alpha)$ and $\widehat{U}_{a}(\Gamma_{a};1-\alpha)$ respectively. These quantiles are used to construct a $(1-\alpha)$ confidence interval (CI), denoted as $\widehat{\text{CI}}_{\textnormal{OR,a}}(\Gamma_{a};1-\alpha)=\left[\widehat{L}_{a}(\Gamma_{a};1-\alpha),\widehat{U}_{a}(\Gamma_{a};1-\alpha)\right]$.
Suppose $\rho_{a}({\bf v})$ is indexed by a finite-dimensional parameter $\boldsymbol{\eta}_{a}$. Theorem 4.1 shows that under regularity conditions, the plug-in, OR estimator $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})$ in equation (4.1) is consistent and the nonparametric, percentile bootstrap leads to a valid CI.
@@ 4.2 Efficient influence function
The analysis pipeline in this section is based on the efficient influence function (EIF) and is more broadly applicable than that in Section 4.1, especially when ${\cal X}$ is not discrete and the propensity score is unknown. However, the EIF-based estimator is more complex and requires estimating multiple nuisance functions.
To motivate the estimator, we first derive the EIF of $\theta_{a}(\Gamma_{a})$ in Theorem 4.2.
@@
$\textnormal{EIF}({\bf O}_{i},\theta_{a}(\Gamma_{a}))=$ $\dfrac{S_{i}w({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\Gamma_{a}}{[\Gamma_{a}\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\pi({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\pi({\bf X}_{i})}\right\}\{Y_{i}-\mu_{a}({\bf X}_{i})\}\right.$
$+\mu_{a}({\bf X}_{i})-\rho_{a}({\bf V}_{i})\bigg{]}+\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\left[\dfrac{\Gamma_{a}\rho_{a}({\bf V}_{i})}{\Gamma_{a}\rho_{a}({\bf V})+1-\rho_{a}({\bf V}_{i})}-\theta_{a}(\Gamma_{a})\right],$
@@
We remark that when transportability holds, i.e., $\Gamma_{a}=1$, Theorem 4.2 reduces to the EIF in Zeng et al., (2023).
Following the modern trend in causal inference, we use cross-fitting and the EIF (e.g., Chernozhukov et al., (2017); Kennedy, (2022)) to estimate $\theta_{a}(\Gamma_{a})$. Specifically, we randomly partition the source and target sample indices ${\cal I}_{s}$ and ${\cal I}_{t}$ into $K$ disjoint sets, ${\cal I}_{s}^{(k)}$ and ${\cal I}_{t}^{(k)}$, respectively, for $k=1,2,\cdots,K$, and let ${\cal I}^{(k)}={\cal I}_{s}^{(k)}\cup{\cal I}_{t}^{(k)}$. For each $k$, the nuisance functions are estimated with data in ${\cal I}\setminus{\cal I}^{(k)}$ and they are denoted as $\widehat{\pi}^{(k)}({\bf x})$, $\widehat{\mu}_{a}^{(k)}({\bf x})$, $\widehat{w}^{(k)}({\bf v})$ and $\widehat{\rho}_{a}^{(k)}({\bf v})$. We then plug them into the “uncentered” EIF and evaluate it with the data in ${\cal I}^{(k)}$:
$\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\Gamma_{a})=$ $\dfrac{1}{|{\cal I}_{s}^{(k)}|}\sum_{i\in{\cal I}_{s}^{(k)}}\dfrac{\Gamma_{a}\widehat{w}^{(k)}({\bf V}_{i})}{[\Gamma_{a}\widehat{\rho}_{a}^{(k)}({\bf V}_{i})+1-\widehat{\rho}_{a}^{(k)}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\widehat{\pi}^{(k)}({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\widehat{\pi}^{(k)}({\bf X}_{i})}\right\}\{Y_{i}-\widehat{\mu}_{a}^{(k)}({\bf X}_{i})\}\right.$
$\left.\quad{}\quad{}\quad{}\quad{}+\widehat{\mu}_{a}^{(k)}({\bf X}_{i})-\widehat{\rho}_{a}^{(k)}({\bf V}_{i})\right]+\dfrac{1}{|{\cal I}_{t}^{(k)}|}\sum_{i\in{\cal I}_{t}^{(k)}}\dfrac{\Gamma_{a}\widehat{\rho}_{a}^{(k)}({\bf V}_{i})}{\Gamma_{a}\widehat{\rho}_{a}^{(k)}({\bf V}_{i})+1-\widehat{\rho}_{a}^{(k)}({\bf V}_{i})}.$
Finally, we take an average of $\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\Gamma_{a})$ to arrive at the EIF-based, cross-fitting estimator of $\theta_{a}(\Gamma_{a})$, which we denote as $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})=K^{-1}\sum_{k=1}^{K}\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\Gamma_{a})$. A step-by-step algorithm can be found from Section C of the Appendix. Theorem 4.3 shows that under conditions, $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$ is consistent, asymptotically normal, and semiparametrically efficient.